# Video In-context Learning

### v1

@@ Abstract
In-context learning for vision data has been underexplored compared with that in natural language. Previous works studied image in-context learning, urging models to generate a single image guided by demonstrations. In this paper, we propose and study video in-context learning, where the model starts from an existing video clip and generates diverse potential future sequences, each semantically guided by the prompted video demonstrations. To achieve this, we provide a clear definition of the task, and train an autoregressive Transformer on video datasets. We thoroughly analyze the effect of different datasets and represent frames as discrete tokens, and then model them by next token predictions. We design various evaluation metrics, including both objective and subjective measures, to demonstrate the visual quality and semantic accuracy of generation results. Our model follows the scaling law and generates high-quality video clips that accurately align with the semantic guidance provided by in-context examples.
Figure 1: Illustrations of video in-context learning. Given a query video clip, the model generates different results based on the provided demonstration video clips. Text descriptions are annotated for clarity and not used as model input. The generated results are coherent with the query in content and semantically consistent with the demonstrations.
@@ 1 Introduction
Large language models (LLMs) [29, 6, 34, 8] demonstrate fascinating abilities to generate different responses to a given query when providing various demonstrations, this in-context learning [6] capacity helps LLMs generalize to zero-shot tasks smoothly. Compared to natural languages, visual demonstrations such as images and videos are abundant and can convey broader information about the physical world (e.g., spatial and temporal information in different scales of resolutions) [42], making the visual in-context learning an appealing and essential capacity for an ideal world model that perceives visual signals and makes responses correspondingly.
Previous works [3, 38, 39, 1] have studied image in-context learning (Img-ICL). By formulating image perception tasks such as segmentation and detection to image pair demonstrations in delicately designed structures, e.g., a joint grid image [3] for MAE models [15] or a sequence [1] for Transformer models [36], they encourage the model to mimic the task and generate the prediction of a given image query. Img-ICL focuses on perception tasks and falls short of generating semantic and sequential responses, while large language models have shown remarkable capabilities in generating coherent and contextually appropriate textual sequences. This gap highlights the potential for extending in-context learning to video data, where the temporal dimension provides dynamic causality and richness that static images cannot provide.
In this paper, we bridge the gap by introducing video in-context learning (Vid-ICL). Unlike image-based approaches, Vid-ICL uses videos as the fundamental unit for both demonstrations and outputs. Demonstration videos are highly versatile and capable of conveying a wide range of information, such as examples for various tasks including moving or grabbing objects, or movements of the camera in an ego-centric video. Guided by these demonstrations, the model receives a query video clip and is expected to generate a subsequent video sequence that is semantically aligned with the demonstrations (see Figure 1 for examples). This allows Vid-ICL to address multiple downstream tasks, such as embodied planning and simulating, by letting a query robot imitate the actions demonstrated by other robots. As videos are good at describing low-level details (where language may fall short) and temporal dynamics (where images are insufficient), Vid-ICL serves as a crucial interface for models to interact with the real world.
@@
We show that an autoregressive Transformer [36, 34] trained with next token prediction demonstrates zero-shot capacity of Vid-ICL, with a thorough analysis and selection on public video datasets. Each training sample comprises a sequence of frames extracted from a single video, encoded as discrete tokens using a vector-quantized encoder [35, 28]. The model learns to predict future frames based on preceding ones through self-supervised learning. Though in-context examples are not provided during training, the autoregressive paradigm allows the model to learn the underlying structures and patterns present in the videos. Consequently, when exposed to in-context examples during inference, the model can generalize and apply these learned patterns to generate coherent and contextually appropriate video sequences. This zero-shot capability mirrors the findings observed in large language models [6].
@@
To comprehensively and accurately evaluate the model performance in Vid-ICL, we develop both objective and subjective metrics to assess the generated videos in terms of visual quality, semantic accuracy, and consistency with the prompted demonstrations. Our extensive experiments demonstrate that the model not only produces high-quality video clips but also successfully adheres to the semantic guidance provided by the in-context examples. In addition, we show that the zero-shot Vid-ICL capacity also follows the scaling law [23] of large models, illustrating the potential of future works.
The main contributions of this work are summarized as follows:
@@
  • We propose and study the task of video in-context learning, which enables the model to interact with real-world demonstrations through video generation.
@@
  • We train a large Transformer model that exhibits powerful video in-context learning capacity, which also follows the scaling law of large models.
@@
  • We propose various evaluation metrics to evaluate the visual quality and semantic accuracy of generated videos, providing a solid benchmark for the evaluation of video in-context learning.
@@ 2 Related work
In-context Learning for Natural Language
In-context learning (ICL) has emerged as a pivotal feature of large language models (LLMs), obviating the need for parameter updates and thereby benefiting the downstream usage of LLMs [6, 34]. A standard practice involves providing several demonstrations and a query, then tasking the LLM with producing the answer. Different prompting designs [40, 14, 25] empower LLMs to deal with a wide range of natural language understanding and generation tasks.
In-context Learning for Vision
Pioneering works construct vision ICL as an image inpainting [3, 38, 39] task. Given multiple query-answer image pairs arranged in a grid image, models are optimized to reconstruct the masked answers under the MAE projective [15]. Recently, LVM [1] flatten image pairs into a sequence and train an auto-regressive Transformer with next token prediction. These models show powerful image in-context learning capacity on perception tasks such as semantic segmentation, object detection, and depth estimation, by identifying the prompting task from demonstration pairs and generating the prediction of the query image. When extending the basic elements (i.e., demonstrations, queries and predictions) from images to videos, unfortunately, existing models struggle to handle the increased complexity as they are not designed to capture the spatiotemporal relationships of inputs. Specifically, though LVM [1] is able to generate consecutive frames of a query video clip, the in-context capacity (e.g., generating different consequences with different demonstrations) is not demonstrated.
@@
The studied Vid-ICL can be considered as a conditional video generation problem, which has recently become a prominent focus in research. Text is the most commonly utilized condition, and various text-to-video models have shown promising results in generating high-fidelity videos [19, 18, 5, 21, 37, 24, 44]. Recently, video generation models are spread to other related domains such as embodied AI [42], where they are utilized in visual planning [9] and simulation [41] with actions and observations as conditions, illustrating the potential of this area. As for model architectures, Diffusion models [17, 27] and Transformers [36] are both widely adopted in different scenarios, and we choose the autoregressive Transformer as the backbone model in this paper, as its abilities to model long-range dependencies and contextual relationships are crucial for in-context learning.
3 Video In-context Learning
We introduce Video In-Context Learning (Vid-ICL) in this section. We begin with the problem definition of both image and video in-context learning in Section 3.1, highlighting the differences between them. We then elaborate on the training and inference pipeline of Vid-ICL in Section 3.2. Dataset selection is discussed in Section 3.3.
Figure 2: The framework of Vid-ICL. Left: Training the model. The data used for training are continuous video clips and the Transformer is trained by next token prediction objective. Right: Video In-context Inference. The model is conditioned on demonstration videos and generates the subsequent frames of a given query video.
@@ 3.1 Problem Definition
Image In-context Learning
Typically, Image In-Context Learning (Img-ICL) follows the setting akin to question answering: the demonstration is composed of $k$ image pairs $D^{I}=\{s^{q}_{i},s^{a}_{i}\}_{i=1}^{k}$, where $s^{q}_{i}\in\mathbb{R}^{(3,n_{r},n_{r})}$ denotes the input image with resolution $n_{r}$ and $s^{a}_{i}$ denotes its target (usually an image with task-specific annotations [1]). Given a query image $x^{I}$, the model predicts its target $y^{I}$ conditioned on the demonstration:
@@
where $f_{\theta}(\cdot)$ represents the utilized vision model such as an MAE-pretrained ViT [38, 3] or an auto-regressive Transformer [1], which is expected to discern the inherent task from image pairs $(s^{q}_{i},s^{a}_{i})$ and process $x$ accordingly.
Video In-context Learning
In contrast to Img-ICL, Vid-ICL takes videos as the basic element. Specifically, the demonstration $D^{V}=\{(s_{i}^{1},\cdots,s_{i}^{n_{i}})\}_{i=1}^{k}$ comprises $k$ video clips, and each clip consists of $n_{i}$ frames. Given a query video clip $x^{V}=(s_{q}^{1},\cdots,s_{q}^{n_{q}})$, the objective of Vid-ICL can be written as:
@@ 3.2 Approach
Training a Transformer for vision tasks typically consists of two stages, 1) training a visual tokenizer such as VQ-VAE [35, 10] to convert each image to discrete tokens; 2) each training sample is constructed as a sequence of tokens to train the Transformer decoder. For the first stage, we utilize a public pretrained checkpoint [28] with 16×16$\times$16 × spatial compression rate as the VQ tokenizer. Our framework also fits for recent tokenizers with both spatial and temporal compressions [43, 44], and we leave it for future work. Formally, for a training video clip $x^{V}=(s^{1},\cdots,s^{n})$ with $n$ frames, the VQ tokenizer converts it to a flat sequence $x^{t}=(t_{1},\cdots,t_{n_{t}},t_{n_{t}+1},\cdots,t_{N})$, where $n_{t}$ denotes the number of quantized ids that represents one frame, and $N=n\cdot n_{t}$ denotes the total length of the sequence. We append special tokens [bos] and [eos] to the front and end of the sequence, and no special tokens are inserted between sequences of different frames.
Then, in the second stage, we follow the architecture of LLaMA [34] utilizing RMSNorm normalizing [45] and Rotary Embeddings [31], and train a Transformer decoder in an autoregressive way:
@@
In-context Inference
The in-context inference format differs from that in the training. Following Equ (2), a number of video clips are selected as the demonstration $D^{V}$, which is appended in front of the query clip $x^{t}_{<j}=(t_{1},\cdots,t_{j-1})$ to let the model generate corresponding responses:
@@
The training and inference processes of the model differ in that the demonstration $D^{V}$ is not provided during training. Despite this, the model exhibits zero-shot Vid-ICL capabilities. This can be attributed to two key reasons. Firstly, no special separation token is inserted between frames, allowing the preceding sequences $t_{<i}$ in Equation (3) to be implicitly viewed in a demonstration-query format, i.e., $P(t_{i}|t_{<i,>j},t_{\leq j})$ where $j<i$. This implies that the model has inherently learned to handle sequences resembling the demonstration-query structure during training. Secondly, the autoregressive nature of the Transformer enables it to seamlessly extend its sequence prediction capabilities to scenarios where the demonstration and query come from different videos, thus facilitating smooth generalization to the in-context learning paradigm. Our experiments further show that training the model with explicit in-context examples does not yield significant improvements over the zero-shot approach. Refer to Sec 5.1 for more discussions.
@@
We mainly focus on videos as the demonstration in this paper, but our approach can be easily extended to acquiring other modalities such as text. To do so, we just need to transfer original text descriptions into latent representations denoted as $c$ by a pretrained language model [30], then take $c$ as an additional condition while training the Transformer (Equ (3)) as well as in-context inference (Equ (4)). We show in experiments Sec 5.4 that our model understands and reacts to both text and video demonstrations.
@@ 3.3 Data
While large language models are trained on vast amounts of data (usually trillions of tokens [20, 34]), high-quality video data is constrained. In addition, Vid-ICL prefers videos that exhibit not only rich content but also clear causal relationships and interactivity. As a result, among various public video datasets, we focus on these accomplish embodied tasks and select two primary datasets as our main training data sources: 1) Ego4d [12], an egocentric video dataset featuring abundant first-person activities; and 2) Kinetics-600 [7], a comprehensive video dataset comprising diverse human activities. Additionally, we incorporate Webvid [2] that contains a large amount of internet videos, to augment the variety of video content. Notably, due to the typically ambiguous and inconsistent semantic information conveyed by internet videos, we only utilize a small fraction of Webvid, and show that simply increasing the data scale by adding more internet videos does not help improve the in-context capacity of the model (refer to Sec 5.4).
To validate the Vid-ICL capability, we choose the Something-Something v2 (SSv2) dataset [11] where each video depicts a basic action that occurs in the physical world, such as moving objects in a direction. These videos convey strong semantic information, which can be utilized to construct in-context demonstrations. We utilize the evaluation split of SSv2 as the evaluation set for all experiments.

### v2

@@ Abstract
People interact with the real-world largely dependent on visual signal, which are ubiquitous and illustrate detailed demonstrations. In this paper, we explore utilizing visual signals as a new interface for models to interact with the environment. Specifically, we choose videos as a representative visual signal. And by training autoregressive Transformers on video datasets in a self-supervised objective, we find that the model emerges a zero-shot capability to infer the semantics from a demonstration video, and imitate the semantics to an unseen scenario. This allows the models to perform unseen tasks by watching the demonstration video in an in-context manner, without further fine-tuning. To validate the imitation capacity, we design various evaluation metrics including both objective and subjective measures. The results show that our models can generate high-quality video clips that accurately align with the semantic guidance provided by the demonstration videos, and we also show that the imitation capacity follows the scaling law. Code and models have been open-sourced.
Figure 1: Illustrations of VidIT. Given a query video clip, the model generates different results based on the provided demonstration video clips. The generated results are coherent with the query in content and semantically consistent with the demonstrations. Text descriptions are annotated for clarity and not used as model input.
@@ 1 Introduction
Humans acquire skills through imitation, a fundamental aspect of learning demonstrated even in early childhood. For instance, when children observe how to hold a spoon, they not only learn this specific action but can also generalize it to grasp different spoons in varied contexts. These demonstrations are primarily conveyed through visual signals, which are abundant in our environment and provide intricate guidance. Similarly, for the development of general AI agents, the ability to learn from visual demonstrations and generalize to new situations is crucial and desirable. This capability to understand and imitate visual signals is not only appealing but also essential for advancing AI towards more human-like learning and interaction.
Previous works (Bar et al., 2022; Wang et al., 2023a; b; Bai et al., 2023) have studied image in-context learning in image-based visual tasks. By formulating image perception tasks such as segmentation and detection to supervised image pair demonstrations in delicately designed structures, e.g., a joint grid image (Bar et al., 2022) for MAE models (He et al., 2022) or a sequence (Bai et al., 2023) for Transformer models (Vaswani et al., 2017), they encourage the model to mimic the task and generate the prediction of a given image query. However, these works require explicitly formulating the training samples in source-target pairs, which restricts in-context inference to a supervised capacity. Furthermore, image-level generation struggles to produce coherent and contextually appropriate sequences.
@@
In this paper, we introduce the Video ImiTator (VidIT) model, an autoregressive Transformer (Vaswani et al., 2017; Touvron et al., 2023) trained with the next token prediction objective on video datasets. VidIT uses videos as the primary unit for both inputs and outputs, with each training sample consisting of a sequence of frames from a single video, encoded as discrete tokens using a vector-quantized encoder (Van Den Oord et al., 2017; Patil et al., 2024). The model learns to predict future frames based on preceding ones through self-supervised learning. Though demonstration-generation pairs are not provided during training, the autoregressive paradigm enables the model to discern the underlying structures and patterns within the videos, leading to the emergence of a zero-shot imitation capability. Consequently, when the model encounters demonstration videos during inference, it can generalize and apply these learned patterns to generate coherent and contextually appropriate video sequences. This zero-shot capability echoes findings observed in large language models (Brown et al., 2020).
Specifically, demonstration videos are highly versatile and capable of conveying a wide range of information, such as examples for tasks including object manipulation, or movements of the camera in an ego-centric video. Guided by these demonstrations, the model receives a query video clip set in a new scene and generates a subsequent video sequence that mimics the semantic actions from the demonstrations (see Figure 1 for examples). This allows VidIT to address multiple downstream tasks, such as embodied planning and simulating, by letting a query robot imitate the actions demonstrated by humans or other agents. Given that videos excel in describing low-level details (where language may fall short) and temporal dynamics (where images are insufficient), VidIT acts as a crucial interface for models to interact with the real world.
@@
To comprehensively and accurately evaluate the model performance in VidIT, we develop both objective and subjective metrics to assess the generated videos in terms of visual quality, semantic accuracy, and consistency with the prompted demonstrations. Our extensive experiments demonstrate that the model not only produces high-quality video clips but also successfully adheres to the semantic guidance provided by the demonstration examples. In addition, we show that the zero-shot imitation capacity also follows the scaling law (Kaplan et al., 2020) of large models, illustrating the potential of future works. The main contributions of this work are summarized as follows:
@@
  • We propose and study the task of video imitation, which enables the model to interact with real-world demonstrations through video generation.
@@
  • We train a large Transformer model VidIT that exhibits powerful video imitation learning capacity, which also follows the scaling law of large models.
@@
  • We propose various evaluation metrics to evaluate the visual quality and semantic accuracy of generated videos, providing a solid benchmark for the evaluation of video imitation learning.
@@ 2 Related work
Imitation Learning and In-context Learning
Imitation Learning is a crucial paradigm in reinforcement learning, wherein the expected behavior is acquired by mimicking demonstration’s actions (Zare et al., 2024). By learning on state-action pairs from an expert, the model is expected to map the current state to the corresponding actions. Such learning paradigm is close to human behavior and is appealing in interactive circumstances.
This ability to imitate has also been identified as a key feature in large language models (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023; Chowdhery et al., 2023), which is referred to as in-context learning in LLM area. By conditioning on a sequence of text templates, an LLM generates coherent content based on these templates. This paradigm obviates the need for parameter updates and thereby benefiting the downstream usage of LLMs (Brown et al., 2020; Touvron et al., 2023).
Different prompting designs (Wei et al., 2022; Guo et al., 2023; Li et al., 2023) empower LLMs to deal with a wide range of natural language understanding and generation tasks.
As for the in-context paradigm in vision area, current research mainly focus on teaching the model to mimic the vision task provided by demonstrations. Pioneering works construct vision imitation as an image inpainting (Bar et al., 2022; Wang et al., 2023a; b) task. Given multiple query-answer image pairs arranged in a grid image, models are optimized to reconstruct the masked answers under the MAE projective (He et al., 2022). Recently, LVM (Bai et al., 2023) flatten image pairs into a sequence and train an auto-regressive Transformer with next token prediction. These models show powerful capacity to imitating vision tasks such as semantics segmentation and detection from the demonstrations, but rely on training on supervised datasets. In addition, when extending the basic elements (i.e., demonstrations, queries and predictions) from images to videos, unfortunately, existing models struggle to handle the increased complexity as they are not designed to capture the spatial-temporal relationships of inputs. Specifically, though LVM (Bai et al., 2023) is able to generate consecutive frames of a query video clip, the video imitation capacity (e.g., generating different consequences with different demonstrations) is not demonstrated.
@@
The studied video imitation learning can be considered as a conditional video generation problem, which has recently become a prominent focus in research. Text is the most commonly utilized condition, and various text-to-video models have shown promising results in generating high-fidelity videos (Ho et al., 2022b; a; Brooks et al., 2024; Hong et al., 2023; Villegas et al., 2023; Kondratyuk et al., 2023; Yu et al., 2023b). Recently, video generation models are spread to other related domains such as embodied AI (Yang et al., 2024), where they are utilized in visual planning (Du et al., 2023) and simulation (Yang et al., 2023) with actions and observations as conditions, illustrating the potential of this area. As for model architectures, Diffusion models (Ho et al., 2020; Nichol & Dhariwal, 2021) and Transformers (Vaswani et al., 2017) are both widely adopted in different scenarios, and we choose the autoregressive Transformer as the backbone model in this paper, as its abilities to model long-range dependencies and contextual relationships are crucial for video imitators.
3 Video Imitator
We introduce the proposed Video Imitator in this section. We begin with the problem definition of both image and video imitation in Section 3.1, highlighting the differences between them. We then elaborate on the training and inference pipeline of VidIT in Section 3.2. Dataset selection is discussed in Section 3.3.
Figure 2: The framework of VidIT. Left: Training the VidIT. The data used for training are continuous video clips and the Transformer is trained by next token prediction objective. Right: Video Imitation Inference. The model is conditioned on demonstration videos and generates the subsequent frames of a given query video.
@@ 3.1 Problem Definition
Typically, the image task imitator models follow the setting akin to question answering: the demonstration is composed of $k$ image pairs $D^{I}=\{s^{q}_{i},s^{a}_{i}\}_{i=1}^{k}$, where $s^{q}_{i}\in\mathbb{R}^{(3,n_{r},n_{r})}$ denotes the input image with resolution $n_{r}$ and $s^{a}_{i}$ denotes its target (usually an image with task-specific annotations (Bai et al., 2023)). Given a query image $x^{I}$, the model predicts its target $y^{I}$ conditioned on the demonstration:
@@
where $f_{\theta}(\cdot)$ represents the utilized vision model such as an MAE-pretrained ViT (Wang et al., 2023a; Bar et al., 2022) or an auto-regressive Transformer (Bai et al., 2023), which is expected to discern the inherent task from image pairs $(s^{q}_{i},s^{a}_{i})$ and process $x$ accordingly.
Unlike image task imitators, VidIT takes videos as basic element, focusing on inheriting semantics from demonstrations rather than discerning specific tasks. Specifically, the demonstration $D^{V}=\{(s_{i}^{1},\cdots,s_{i}^{n_{i}})\}_{i=1}^{k}$ comprises $k$ video clips, and each clip consists of $n_{i}$ frames. Given a query video clip $x^{V}=(s_{q}^{1},\cdots,s_{q}^{n_{q}})$, the objective of VidIT can be formulated as:
@@ 3.2 Approach
Training a Transformer for vision tasks typically consists of two stages, 1) training a visual tokenizer such as VQ-VAE (Van Den Oord et al., 2017; Esser et al., 2021) to convert each image to discrete tokens; 2) each training sample is constructed as a sequence of tokens to train the Transformer decoder. For the first stage, we utilize a public pretrained checkpoint (Patil et al., 2024) with 16×16$\times$16 × spatial compression rate as the VQ tokenizer. Our framework also fits for recent tokenizers with both spatial and temporal compressions (Yu et al., 2023a; b), and we leave it for future work. Formally, for a training video clip $x^{V}=(s^{1},\cdots,s^{n})$ with $n$ frames, the VQ tokenizer converts it to a flat sequence $x^{t}=(t_{1},\cdots,t_{n_{t}},t_{n_{t}+1},\cdots,t_{N})$, where $n_{t}$ denotes the number of quantized ids that represents one frame, and $N=n\cdot n_{t}$ denotes the total length of the sequence. We append special tokens [bos] and [eos] to the front and end of the sequence, and no special tokens are inserted between sequences of different frames.
Then, in the second stage, we follow the architecture of LLaMA (Touvron et al., 2023) utilizing RMSNorm normalizing (Zhang & Sennrich, 2019) and Rotary Embeddings (Su et al., 2024), and train a Transformer decoder in an autoregressive way:
@@
Imitation Inference
The imitation inference format differs from that in the training. Following Equation (2), a number of video clips are selected as the demonstration $D^{V}$, which is appended in front of the query clip $x^{t}_{<j}=(t_{1},\cdots,t_{j-1})$ to let the model generate corresponding responses:
@@
The training and inference processes of the model differ in that the demonstration $D^{V}$ is not provided during training. Despite this, the model exhibits zero-shot video imitation capabilities. This can be attributed to two key reasons. Firstly, no special separation token is inserted between frames, allowing the preceding sequences $t_{<i}$ in Equation (3) to be implicitly viewed in a demonstration-query format, i.e., $P(t_{i}|t_{<i,>j},t_{\leq j})$ where $j<i$. This implies that the model has inherently learned to handle sequences resembling the demonstration-query structure during training. Secondly, the autoregressive nature of the Transformer enables it to seamlessly extend its sequence prediction capabilities to scenarios where the demonstration and query come from different videos, thus facilitating smooth generalization to the imitation paradigm. Our experiments further show that training the model with explicit imitation examples does not yield significant improvements over the zero-shot approach. Refer to Sec 5.1 for more discussions.
@@
We mainly focus on videos as the demonstration in this paper, but our approach can be easily extended to acquiring other modalities such as text. To do so, we just need to transfer original text descriptions into latent representations denoted as $c$ by a pretrained language model (Raffel et al., 2020), then take $c$ as an additional condition while training the Transformer (Equation (3)) as well as imitation inference (Equation (4)). We show in experiments Sec 5.5 that our model understands and reacts to both text and video demonstrations.
@@ 3.3 Data
While large language models are trained on vast amounts of data (usually trillions of tokens (Hoffmann et al., 2022; Touvron et al., 2023)), high-quality video data is constrained. In addition, VidIT prefers videos that exhibit not only rich content but also clear causal relationships and interactivity. As a result, among various public video datasets, we focus on these accomplish embodied tasks and select two primary datasets as our main training data sources: 1) Ego4d (Grauman et al., 2022), an egocentric video dataset featuring abundant first-person activities; and 2) Kinetics-600 (Carreira et al., 2018), a comprehensive video dataset comprising diverse human activities.
Additionally, we incorporate self-collected videos that contains a large amount of general real-world videos, to augment the variety of video content.
To validate the VidIT’s imitation capability, we choose the Something-Something v2 (SSv2) as the main evaluation dataset (Goyal et al., 2017) where each video depicts a basic action that occurs in the physical world, such as moving objects in a direction. These videos convey strong semantic information, which can be utilized to construct the demonstrations. We utilize the evaluation split of SSv2 as the evaluation set for all experiments. In addition, we include the Robotics Transformer-1 (RT-1) (Brohan et al., 2022) dataset and MineRL [^1] https://github.com/minerllabs/minerl to demonstrates VidIT’s effectiveness on embodied AI and interactive tasks.